# Néel-Vector-Orientation Induced Intrinsic Half-Metallicity in Two-Dimensional Altermagnets

### v1

@@ Acknowledgements.
We would like to thank Shandong Institute of Advanced Technology and National Supercomputing Center (Shuguang) for providing computational resources. This work is supported by the National Natural Science Foundation of Shandong Province(Grant No. ZR2024QA040). Xin Chen thanks China scholarship council for financial support (No. 201606220031). Duo Wang acknowledges financial support from the Science and Technology Development Fund from Macau SAR (Grant No. 0062/2023/ITP2 and No. 0016/2025/RIA1) and the Macao Polytechnic University (Grant No. RP/FCA-03/2023). Biplab Sanyal acknowledges financial support from Swedish Research Council (grant no. 2022-04309), STINT Mobility Grant for Internationalization (grant no. MG2022-9386) and DST-SPARC, India (Ref. No. SPARC/2019-2020/P1879/SL).

### v2

@@ Acknowledgements.
We would like to thank Shandong Institute of Advanced Technology and National Supercomputing Center (Shuguang) for providing computational resources. This work is supported by the National Natural Science Foundation of Shandong Province(Grant No. ZR2024QA040). Xin Chen thanks China scholarship council for financial support (No. 201606220031). Duo Wang acknowledges financial support from the Science and Technology Development Fund from Macau SAR (Grant No. 0062/2023/ITP2 and No. 0016/2025/RIA1) and the Macao Polytechnic University (Grant No. RP/FCA-03/2023). Biplab Sanyal acknowledges financial support from Swedish Research Council (grant no. 2022-04309) and STINT Mobility Grant for Internationalization (grant no. MG2022-9386).